OpenStudy (anonymous):

Solve the equation. 6b + 4 = 2(2b + 1)

OpenStudy (anonymous):

6b + 4 = 4b +2 2b + 4 = 2 2b = -2 b = -1

OpenStudy (anonymous):

you use the distributive property- 6b+4=2(2b)+2(1) then you put together like terms and add